A Complete Bibliography of Publications in Theory of Probability and its Applications: 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 October 2023 Version 3.03

Title word cross-reference

d [SS20]. f [LCY^{+}22]. F^{α} [HB20]. G [PY22]. k [PS21a]. m [MQY22]. M|G|1 [Afa21]. q [LG20]. r [GCV20]. ρ [LYJ^{+}22]. U [Gir21].

-Correlograms [GCV20]. -Dependent [MQY22]. -Expectation [PY22]. -Functional [LG20]. -Guaranteed [SS20]. -Mixing [LYJ^{+}22]. -Moment [LCY^{+}22]. -Schemes [HB20]. -Statistics [Gir21]. -th [PS21a].

2021 [Yar22b]. 2022 [Yar22c, Yar23].

4th [SP20].

75th [Edi21a]. 7th [SP23a, SP23b].

90th [CBTV23, Edi21b, Kha23].

A. [Edi21b, Edi21a]. Abstracts [SP20, SP23a, SP23b]. Accompanying [PS22]. Accuracy [Len23, Roo22]. Added [CD20]. Afanasyeva [Edi23]. Allocation [MQY22, Rok21]. Alternative [GZ22, Nag21a]. Analogue [Pla22]. Application [LCY^{+}22, LWW21]. Applications [FKL20b, FS22, LYJ^{+}22, PS21b]. Approach [KST22, MGD22, Smi23]. Approximate [NP20]. Approximating [GZ22]. Approximation [CG22, ISF21, PT21, SX21]. Arbitrage [BS20]. Arcsine [BS23]. Arrays [XLCW23]. Asian [MPP23]. Asset [CKT20, SV20]. Assignment [LT23]. Associated [ISF22]. Association [Tu21]. Asymptotic [AB23, BS20, DN21, Kha22, Pin23, PS22, SK22, SZ21]. Asymptotics [AK22, BMP20, Mos21, PW22, PR23, Roz23, SS20]. Attaining [Afa20, Mos21]. Atypical [VD23]. Autoregression [Bol20].

B. [PS22]. Background [Jur22]. Backward [ABMP21, BT22, Bel21]. Balls [LMW23]. Banking [EMP21]. Based [MS23, MQY22, Pal20]. Be [YSK20]. Behavioral [CR20]. Bernoulli [BN22, BM20]. Bernstein [RS21]. Berry [LYJ^{+}22]. Bertrand [Kau22]. Bessel [Sav21]. Between [SZ21, Smi23]. Bicentenary [Shi22b]. Bifractional [LMW23]. Binary [DNRJ23]. Birth [Shi22b]. Birthday [CBTV23, Edi21b, Kha23, Shi22a, Edi21a]. Blocks [Bru22]. Blotto [Kha22]. Books [Zhi23]. Borovkov [Edi21b]. Boundary [KL21, KK23, Lee21]. Bounded [PC21]. Bounds [Iva23, LYJ^{+}22, Pin22]. Branching [Bal22, Bru22, Bul20, Leb22, PR23, VD20, VD21, VD23]. Breaking [ABL21]. Brownian [ABL21, AK22, BCE23, Gra20, Lee21, LMW23, Sav20]. Building [Bru22].

Call [MPP23]. Case [Har22, Khv23]. Catalytic [Bul20]. Cauchy [PT21]. Centenary [Edi20c]. Central [BU22, IPF21, Roo22, SK22]. Certain [KF21]. Chain [ABL21]. Chains [Nag21a]. Characteristic [Roo22]. Characterization [CY22, PS21a]. Characterizing [PS21b]. Chebyshev [Bin22, BU22, Bog22, BLM22, Shi22b]. Chebyshev-Type [BLM22]. Checkable [SLK20]. Chi [Bol20, Sav21]. Chi-Square [Bol20, Sav21]. Chibisov [Edi20a]. Chromatic [DS22]. Circle [Har22, SX21]. Class [GH21]. Classes [KF21]. Classical [FKL20b]. Close [Afa21]. Close-Down [Afa21]. CLT [Roz20]. Clustering [Bal22]. Coefficients [Zad21]. Colonel [Kha22].

Combinatorial [Roo22]. Common [BT22]. Communication [Rok21]. Compatible [CPRS20]. Complement [ZL20]. Complete [CY22, LCY^{+}22, LWW21, XLCW23]. Complex [Ale22, Ale23]. Complexity [Dar20]. Components [AB23]. Compound [Bak21, BMP20, Bor21a, Bor21b, Tim21]. Concentration [Sko23]. Condition [FKL20a, Lee21]. Conditional [BBC22, MS21]. Conditions [CD20, IPF21, SLK20, YSK20]. Cone [AH20]. Conference [SP20, SP23a, SP23b]. Conferences [Zhi23]. Configuration [PC21]. Conic [CR20]. Connections [AH20]. Constant [NP20]. Constraints [Smi23]. Construction [MS21]. Continuous [CKT20, Dar20, Gus21, Kab21]. Control [AHØP22, FS22, Pal22]. Controllable [Gra20]. Controlled [BT22]. Convergence [BS23, FKL20b, KF21, LCY^{+}22, LWW21, MGD22, Roz23, Sle21, XLCW23, ZSG21]. Converging [FKL20a]. Convolutions [BM20, CG22]. Cooperative [EMP21]. Copula [MS21]. Correction [BU22]. Correlation [Lin23, Pin23]. Correlograms [GCV20]. Cost [Afa23, Pal22, Pal23b]. Costs [MPP23, NP20]. Covariance [Yas22]. Cox [dRdC20]. Critical [Khv23]. Curve [BSZ23].

D. [Edi20a]. Data [Gir21]. Dawson [Li21]. Decision [FKL20b, Piu23]. Decomposable [Jur22, VD23]. Decreasing [Roz22]. Degree [PC21]. Delaunay [EN20]. Delays [Afa22]. Density [Har22, Lee21]. Department [Yar22b, Yar22c, Yar23]. Dependence [FF20, Tu21]. Dependent [AB23, LCY^{+}22, Lee21, LWW21, MQY22, SDDT22, SZ21, XLCW23, Yas22]. Design [Lin23]. Determinacy [SLK20]. Determined [YSK20]. Deterministic [Smi23]. Deviation [BBC22, BMP20, Bor21b, BLM22, LH23].

Deviations [Bak21, Bor21a, Roz22, dRdC20]. Device [Afa22]. Dickman [MS23]. Differential [BP20, Gli21, Zad21]. Diffusion [Ark20, BT22, Har22]. Diffusions [ABMP21, KST22]. Dimension [Dar20]. Dimensional [Dar20, Gra20, Har22, Sav21]. Dirichlet [Lee21, MS23]. Discounting [BS20]. Discrete [Kov21, Tim21]. Discrimination [SS20]. Distance [MGD22, ZSG21]. Distinguishing [KR21]. Distribution [APY22, AH20, BMP20, Har22, Jur22, KL21, KR21, Mos21, PW22, SX21, YSK20, Yas22, Zor20]. Distributional [PY22]. Distributions [Bog22, GZ22, GH21, MS23, MS21, PC21, Roz22, Sav21, Sle21, SLK20, Tim21]. Divisible [CG22]. Doctor [BDS22]. Domains [Lee21]. Down [Afa21]. Drift [Gra20, Mos21, dRdC20]. Driving [Jur22]. Dual [Rok21]. Duality [CY22].

Edges [PC21]. Edgeworth [BU22]. Editors [Yar22a]. Effect [Bal22]. Efficiency [Pin23]. Eigenvalue [OW23]. Elements [Yas22]. Empirical [SZ21]. Engelbert [BDS22]. Entropy [Kov21]. Environment [Afa20, LH23, VD20, VD21]. Equation [Lee21, PT21, Ver21]. Equations [Bel21, BP20, FS22, Gli21, LG20, Zad21]. Ergodic [Nag21a]. Ergodicities [Li21]. Esseen [LYJ^{+}22]. Estimate [Nag21b]. Estimation [BN22, Len23]. Estimators [BSZ23, Lin23]. Evolution [Tu21, VD20]. Exact [Bor21b, Pin22]. Excitations [Pal23a]. Excursion [PW22]. Excursions [Kre20]. Existence [CY22]. Exit [Har22, KL21]. Expansion [AB23]. Expansions [LMW23, PS22, SK22]. Expectation [AB23, PY22, Zad21]. Expectations [XLCW23, ZSG21]. Expected [Iva23]. Explicit [AD22]. Explosive [dRdC20]. Exponent [AK22]. Exponential [Gir21, Li21]. Expressions [AD22]. Extended [LCY^{+}22, XLCW23]. Extension [Bor21a]. External [Pal23a]. Extrema [MQY22]. Extreme [BBC22]. Extremes [Leb22, MGD22].

Faculty [Yar22b, Yar22c, Yar23]. Failures [FF20]. Fall [Yar23]. Families [ISF22]. Family [VD21]. Fatalov [Ali20]. Fatou [FKL20b, FKL20a]. Feller [Tu21]. Feynman [Pla22]. Field [BT22]. Filtration [CKT20]. Financial [BS20, CKT20, NP20, PY22, Zhi20]. Finite [Dar20, Piu23]. Finite-Dimensional [Dar20]. First [Har22, KK23, Lee21]. First-Passage [KK23]. Flights [KF21]. Flow [AHØP22, BP20, KST22]. Fluctuations [Bul20]. Forest [Khv23]. Form [AH20, FKL20b]. Forms [Tu21]. Formula [KM22, Pla22]. Forward [Bel21]. Fractional [AK22, BCE23, LMW23, Sav20]. Fractionally [AK22]. Free [CG22]. Front [Bul20]. Function [BN22, BMP20, Roo22]. Functional [LG20]. Functionals [ISF21, Pal20]. Functions [AH20, Jur22, Pal20, Sav20, Zor20]. Fundamental [CKT20].

G [Edi23]. Gale [Zhi20]. Galois [AH20]. Galton [Bru22, Khv23]. Game [Kha22]. Games [BT22]. Gamma [Jur22]. Gaussian [BCE23, Bog22, KL21, Kre20, Pin22, Sko23, SX21]. Gaussianity [GCV20]. General [AHØP22, Nag21a, Yar22b, Yar22c, Yar23]. Generalization [Nag21b]. Generalized [Bel21, MS23, MQY22, SDDT22]. Geometric [Sle21]. Gibbs [BBC22]. Given [Khv23, PC21, SP20, SP23a, SP23b]. Global [BSZ23]. Gnedenko [PS22]. Gradient [KST22, Rok21]. Graph [PC21, PR23, Yas22]. Graph-Dependent [Yas22]. Great [Shi22b]. Grigoriev [ZL20]. Growing [KL21]. Guaranteed [SS20, Smi23]. Gumbel [ABL21].

Half [DG20]. Half-Spaces [DG20]. Hans [BDS22]. Hans-Jürgen [BDS22]. Hardy [NP21]. Heat [Lee21]. Heavy [AB23]. Heavy-Tailed [AB23]. Hedging [MPP23, NP20]. Heredity [Leb22]. Hermite [Bog22]. Hermitian [OW23]. High [Afa20, BSZ23, Kre20, Pla22]. High-Order [Pla22]. Higher [PT21]. Higher-Order [PT21]. Hilbert [SDDT22]. Holds [Jan21]. Hua [AD22]. Hypergraph [DS22]. Hypotheses [SS20]. Hypothesis [DNRJ23].

I. [Kha23]. I.I.D. [Gir21, NP21]. Ibragimov [Kha23]. II [Ale23, ISF22, SP23b]. Immigration [VD21]. Incentive [CPRS20]. Incentive-Compatible [CPRS20]. Increases [Leb22]. Independence [DN21]. Independent [BCE23, BU22, Kov21, Roz20, Roz22]. Indetermined [LG20]. Index [Ale22, Ale23]. Indicators [HB20]. Inequalities [BLM22, GNT20, KN22]. Inequality [Gir21]. Infinitely [CG22]. Influential [DG20]. Information [AHØP22, CD20, DNRJ23, Yar22b, Yar22c, Yar23]. Ingersoll [dRdC20]. Inhomogeneous [Har22]. Initial [Har22, VD20]. Insensitivity [Lin23]. Institute [KT23]. Integer [BU22]. Integer-Valued [BU22]. Integrability [BP20, FKL20a, ORÜV22]. Integral [BSZ23, Pal20, XLCW23]. Integrated [AK22]. Integro [Roz20]. Integro-Local [Roz20]. Intensity [Mos21]. Interacting [ABL21]. Interarrival [AB23]. International [SP20, SP23a, SP23b]. Interruptions [Afa22]. Intersections [LL23]. Invariance [Bor21a]. Invariant [LWW21]. Inverse [KK23, Pal23b]. Inversely [Pal22]. Investment [SV20]. Investors [CR20].

Jackson [Len23]. Joint [Gus21, PW22].

Jump [FS22]. Jumps [BP20, KN22, NP20]. Jürgen [BDS22].

Kabanov [ZL20]. Kac [Pla22].

Kantorovich [KF21]. Kendall [Pin23]. Khasanovich [Edi20c]. Known [Iva23]. Kolmogorov [Bel21, FS22, GZ22, Jan21].

L [Bin22, Shi22b]. L. [Edi23]. Langevin [KST22]. Laplace [BMP20, Sle21]. Large [Bak21, BS20, Bor21a, Bor21b, BLM22, CKT20, Jan21, NP21, Rok21, Roz22, Yas22, Zor20, dRdC20]. Last [KL21]. Lattice [APY22]. Law [BS23, Gus21, Jan21, Sle21, TJIN22, NP21]. Laws [CG22, Kab21, SDDT22]. Lebesgue [FKL20b]. Legacy [Bin22]. Lemma [FKL20b, FKL20a]. Length [PW22]. Lengths [Len23]. Letter [Yar22a]. Levels [Afa20]. Lévy [ISF22]. Life [Bin22, Yar22b, Yar22c]. Likelihood [GKKT22]. Limit [ABL21, BM20, BU22, DS22, GZ22, HB20, IPF21, PC21, PS22, Roo22, Sav20, SK22, Tim21, TJIN22]. Limiting [Yas22]. Lindeberg [IPF21]. Linear [GNT20, ISF22, KL21, Leb22, LWW21, Mos21, OW23, Pal22, Pal23b, Pal23a, Zad21]. Linear-Quadratic [Pal23b]. Linear-Time-Invariant [LWW21]. Littlewood [NP21]. Local [BM20, Gus21, ISF21, LMW23, PW22, Roz20, Tim21]. Log [CY22, Jur22]. Log-Gamma [Jur22]. Log-Optimal [CY22]. Lomonosov [Yar22b, Yar22c, Yar23]. Lower [AH20, Pin22].

M [Edi20a, Edi20b]. Mallows [MGD22]. Malyshev [ILM^{+}23]. Maps [Dar20]. Marcinkiewicz [SDDT22]. Marginals [GH21]. Market [CR20, Zhi20]. Markets [BS20, CKT20, NP20]. Markov [FKL20b, FS22, Nag21a, Piu23]. Martingale [BCE23]. Martingales [Kab21]. Mass [Sko23]. Matching [GH21]. Mathematical [AB23]. Mathematician [Shi22b]. Mathematics [Yar22b, Yar22c, Yar23]. Matrices [OW23, Yas22]. Max [Gus21, Leb22]. Max-Continuous [Gus21]. Max-Linear [Leb22]. Maxima [Iva23]. Maximal [KN22, LL23, NP21]. Maximum [Gus21, Kov21, LT23, Mos21, PW22]. McKean [BT22, Ver21]. Mean [BT22]. Measured [PY22]. Measures [FKL20b, FKL20a, Pin22, PS22]. Mechanics [Yar22b, Yar22c, Yar23]. Meixner [GH21]. Memoriam [BDS22]. Memory [Ali20, Edi20a, Edi20b, Edi23, ILM^{+}23, Shi21]. Method [BN22, Nag21a, Rok21, Roo22]. Methods [ORÜV22, SP20, SP23a, SP23b]. Metric [KF21]. Microcredit [EMP21]. Minimum [SS20]. Missing [Sko23]. Mixing [LYJ^{+}22]. Model [EMP21, RS21, ZL20, Zhi20]. Models [BSZ23, CR20]. Moderately [Bor21a]. Modifications [IPF21]. Moment [GNT20, LCY^{+}22, LG20, LWW21, PR23, SLK20]. Moments [PS21a, SK22, YSK20]. Mosaics [EN20]. Moscow [Yar22b, Yar22c, Yar23]. Motion [AK22, BCE23, Gra20, Lee21, Sav20]. Motions [LMW23]. Multidimensional [APY22, Ver21]. Multitype [Bal22]. Multivariate [GZ22, Leb22, MS21]. Mutually [Pal23b].

Nadaraya [Lin23]. Nagaev [CBTV23]. Negatively [LCY^{+}22, XLCW23]. Neighborhood [Har22]. Networks [Len23, Rok21]. Neumann [Zhi20]. News [Yar22b, Yar22c]. NFLVR [CY22]. No [BS20]. Node [TJIN22]. Noise [BT22]. Non [OW23]. Non-Hermitian [OW23]. Nondeterminism [LMW23]. Nonlinear [ABMP21, Bel21, GNT20]. Nonstandard [GKKT22]. Normal [PS21a, TJIN22]. Normality [Bol20]. Note [LYJ^{+}22].

Novikov [Edi21a]. Number

[DS22, Khv23, PC21, Rok21, SS20, Zor20]. Numbers [Jan21, NP21].

Observations [SS20]. Observers [LWW21]. Open [Len23]. Operation [Afa22]. Operator [Pla22]. Operators [ISF22]. Optimal [AHØP22, CY22, DNRJ23, Pal22, Pal23b, Pal23a]. Optimality [Ark20, BSZ23]. Optimization [Afa23]. Options [MPP23]. Order [BSZ23, BT22, PT21, Pla22]. Organized [KT23]. Ornstein [KN22, Pal20]. Osmotic [Gli21]. Other [BM20, Tu21]. Outliers [Bol20].

P. [Bin22, Shi22b]. Pair [MS21]. Pairwise [Jan21, NP21]. Paradox [Kau22]. Parameter [dRdC20]. Partial [OW23]. Particle [Leb22]. Particles [ABL21, PR23, SX21]. Passage [KK23, Lee21]. Path [KM22]. Pathwise [Ver21]. Pearson [Bol20]. Periodic [PR23]. Persistence [AK22]. Pickrell [AD22]. Point [APY22, Har22]. Poisson [EN20, MS23, Mos21]. Polynomial [Pal23a]. Polynomials [Bog22]. Population [PR23, VD23]. Portfolio [CY22]. Positive [Gra20, Tu21]. Posterior [CPRS20]. Power [ORÜV22]. Precise [Roz23]. Preferences [Pal23b]. Prices [Zhi20]. Pricing [CKT20]. Principle [Bor21a]. Principles [Bor21b, BLM22]. Probabilistic [LL23, PT21, RS21, Smi23]. Probabilities [CPRS20]. Probability [SLK20, YSK20, Yar22b, Yar22c, Yar23]. Problem [Gra20, KK23, KM22, MPP23, PT21, Smi23, SV20]. Problems [FS22, LG20]. Process [AB23, Bak21, BMP20, Har22, ISF21, KL21, LT23, Mos21, Pal20, Pal23a, Sav21, VD20, VD23, dRdC20]. Processes [Ark20, Bor21a, Bor21b, Bru22, FKL20b, FS22, GKKT22, ISF22, KN22, Kre20, Leb22, Li21, LWW21, Piu23, SZ21, VD21]. Profile [TJIN22]. Progressively [CD20]. Prokhorov [ZSG21]. Proof [For22, Nag21a]. Propagation [Bul20]. Properties [BMP20, KST22, LL23, SZ21]. Proportional [NP20, Pal22]. Protected [TJIN22].

Quadrant [Gra20]. Quadratic [Pal20, Pal22, Pal23b]. Quantile [MS21, SZ21]. Quantiles [AH20]. Quenched [LH23]. Queue [Afa21, Len23]. Queueing [Afa22, Afa23].

R [Ali20]. Random [Afa20, Ale22, Ale23, APY22, Bal22, Bog22, DS22, ISF21, Iva23, Jan21, Jur22, KF21, Kov21, LCY^{+}22, Leb22, LT23, LYJ^{+}22, LH23, NP21, OW23, PW22, Piu23, PR23, Roz20, Roz22, Sle21, SDDT22, SZ21, TJIN22, VD20, VD21, XLCW23, Zad21]. Randomized [AHØP22]. Randomly [LCY^{+}22]. Rapidly [Roz22]. Rate [BSZ23, BMP20, Roz23, Sle21]. Rates [BS23, ZSG21]. Ratio [GKKT22]. Record [HB20, PS21a]. Records [Leb22]. Recursive [TJIN22]. Reflecting [ISF22]. Regeneration [MGD22]. Regression [BN22]. Regularity [BP20]. Regulator [Pal23b, Pal23a]. Related [LG20]. Relations [PS21a]. Relationship [Smi23]. Relative [Pin23]. Remark [Zor20]. Remarks [DN21]. Renewal [Bak21, BMP20, Bor21a, Bor21b, AB23]. Repelling [SX21]. Representations [Jur22]. Reproducibility [MS21]. Resource [Rok21]. Reward [AB23]. Risky [SV20]. Role [Bru22]. Ross [dRdC20]. Rotar [IPF21]. Rowwise [XLCW23]. Ruin [Gra20, SV20]. Russia [Yar22b, Yar22c, Yar23]. Russian [Shi22b].

Sagdy [Edi20c]. Sakhanenko [For22]. Sample [FF20, Yas22]. Samples [Iva23].

Scheme [EMP21, MQY22]. Schemes [HB20]. Schrödinger [PT21]. Scientific [Yar22b, Yar22c]. Scores [Leb22]. SDEs [BT22]. Second [BT22, GZ22]. Self [Jur22]. Self-Decomposable [Jur22]. Semigroup [AD22]. Semimartingale [SV20]. Seminar [Yar22b, Yar22c, Yar23]. Seminars [Zhi23]. Senatov [Shi21, SK22]. Sequence [MQY22]. Sequential [DNRJ23]. Sergey [CBTV23]. Series [Jur22, LMW23, Nag21b, ORÜV22, PY22]. Set [AH20]. Set-Valued [AH20]. Sets [Kab21, LL23]. Setting [CKT20]. Settings [GKKT22]. Shape [KK23, Kre20]. Sharp [dRdC20]. Shiryaev [KK23]. Sided [SS20]. Simplified [MS21]. Single [VD21]. Singular [AHØP22]. Sirazhdinov [Edi20c]. Size [Iva23, Khv23, PR23, VD23]. Slowly [KL21]. Small [Bor21a, LH23]. Smaller [Iva23]. Smoluchowski [KST22]. Smooth [SZ21]. Smoothed [BM20]. Smoothing [ABMP21]. Sock [KM22]. Sojourn [APY22]. Solution [Kau22, PT21, Zad21]. Solutions [Bel21, Ver21]. Solvability [Gli21]. Some [SZ21]. Sorting [KM22]. Spaces [DG20, SDDT22]. Spearman [Pin23]. Special [LT23]. Spectral [Yas22]. Spheres [LMW23]. Spitzer [Nag21b]. Spring [Yar22c]. Square [Bol20, Sav21]. Stability [Ale22, Ale23]. Stable [Ale22, Ale23]. Stage [Sav21, VD20]. State [LWW21, Yar22b, Yar22c, Yar23]. Statements [Smi23]. Stationary [Kre20]. Statistical [LCY^{+}22]. Statistics [Gir21, GNT20, OW23, Pin23, PS21b]. Stochastic [BP20, EMP21, Gli21, Kha22, Pal22, Pal23b, Rok21, SP20, SP23a, SP23b, Zhi20]. Stopping [AHØP22, Ark20]. Strategies [Kha22]. Strong [Jan21, LMW23]. Strongly [SX21]. Structure [CD20]. Student [KT23]. Sub [GCV20, Sko23]. Sub-Gaussian [Sko23]. Sub-Gaussianity [GCV20]. Subcritical [VD20, VD21].

Sublinear [XLCW23, ZSG21]. Submartingale [Gus21]. Subordinator [MS23]. Sufficient [SS20]. Sum [BCE23, Kov21, Nag21b, Roz22, Sle21]. Summability [ORÜV22]. Summands [BU22]. Sums [GH21, LCY^{+}22, Roz20]. Supercritical [Leb22, PR23]. Superhedging [Smi23]. Supporting [Zhi20]. Surveys [CPRS20]. Survival [VD21]. Switch [Mos21]. System [Pal23b, Zad21]. Systems [Afa22, Afa23, Bel21, LWW21, Tu21].

Tail [FF20, PW22]. Tailed [AB23]. Tails [KR21, Roz22]. Talks [SP20, SP23a, SP23b]. Tensor [BSZ23]. Term [Yar22c, Yar23]. Terminating [Bak21]. Test [Bol20, Sav21]. Testing [DNRJ23]. Tests [KR21]. th [PS21a]. Their [Bru22, FS22, LL23, LYJ^{+}22, PS21b]. Theorem [ABL21, BBC22, CKT20, For22, Nag21a, NP21, ZL20, BU22, GZ22, IPF21, PS22, Roo22, SK22]. Theorems [BM20, FKL20b, HB20, Sav20, Tim21]. Threshold [Ark20, HB20]. Time [APY22, CKT20, ISF21, KL21, KK23, Lee21, LWW21, LH23, Mos21, Pal20, Pal22, Pal23b, PY22]. Time-Varying [Pal20]. Time-Weighting [Pal22]. Times [Afa20, Afa21, AB23, Ark20]. Trading [Smi23]. Trajectorial [KST22]. Trajectory [LH23, Mos21]. Transaction [MPP23, NP20]. Transform [BMP20]. Trees [Khv23, TJIN22]. Tributes [Shi22a]. Turnpikes [Piu23]. Two [CKT20, DS22, Gra20, Har22, Iva23, SS20, Sav21, VD23]. Two-Dimensional [Gra20, Har22, Sav21]. Two-Sided [SS20]. Two-Stage [Sav21]. Two-Type [VD23]. Type [BLM22, Li21, NP21, VD23].

Uhlenbeck [KN22, Pal20]. Uncertainty [PY22]. Uncorrelated [Jan21]. Uniform [FKL20a, GZ22, ORÜV22]. Uniquely [YSK20]. Uniqueness [KK23, Ver21]. Unit [SX21]. University [Yar22b, Yar22c, Yar23]. Upper [Pal20, Pin22]. Usage [DNRJ23]. Users [Rok21].

V [PS22, Shi21]. V.

[Ali20, Edi20b, ILM^{+}23, Shi21]. Vacation [Afa21]. Vacations [Afa22, Afa23]. Valued [AH20, BU22]. Values [DS22, PS21a, Zor20]. Variable [DG20]. Variables [Ale22, Ale23, Bog22, Jan21, Jur22, Kov21, LCY^{+}22, LYJ^{+}22, NP21, Roz22, SZ21, XLCW23, Zor20]. Varying [FKL20b, Pal20]. Vector [Len23]. Vectors [Roz20, SDDT22]. Vega [KT23]. Velocities [Gli21]. Version [GZ22, ORÜV22]. Vertices [PC21, PR23]. via [CPRS20, ORÜV22]. Viktorovich [CBTV23]. Vlasov [BT22, Ver21].

Walk [Afa20, APY22, Bal22, Bul20, LH23, PW22, Piu23, PR23]. Walks [ISF21]. Watanabe [Li21]. Watson [Bru22, Khv23, Lin23]. Weakly [FKL20a, VD20]. Weighted [EN20, LCY^{+}22, SDDT22]. Weighting [Pal22]. Wiener [ISF21]. without [CY22, Smi23]. Work [Bin22]. Workshops [KT23].

Zipf [LL23]. Zolotarev [Edi20b]. Zones [Bor21a].

References

Aliyev:2023:AEM

[AB23] R. Aliyev and V. Bayramov. On asymptotic expansion for mathematical expectation of a Renewal– Reward process with dependent components and heavy-tailed interarrival times. Theory of Probability and its Applications, 67 (4):645–651, ???? 2023. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aurzada:2021:BCI

[ABL21] F. Aurzada, V. Betz, and M. Lifshits. Breaking a chain of interacting Brownian particles: a Gumbel limit theorem. Theory of Probability and its Applications, 66(2):184–208, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anderson:2021:BNS

[ABMP21] B. D. O. Anderson, A. N. Bishop, P. Del Moral, and C. Palmier. Backward nonlinear smoothing diffusions. Theory of Probability and its Applications, 66(2): 245–262, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arista:2022:EEH

[AD22] J. Arista and N. Demni. Explicit expressions of the Hua– Pickrell semigroup. Theory of Probability and its Applications, 67(2):208–228, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T990885.

Afanasyev:2020:TAH

[Afa20] V. I. Afanasyev. On the times of attaining high levels by a random walk in a random environment. Theory of Probability and its Applications, 65(3):359–374, ???? 2020. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasyev:2021:VQC

[Afa21]  G. A. Afanasyev.  A vaca-

tion queue M|G|1 with closedown times. Theory of Probability and its Applications, 66(1):1–14, ???? 2021. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasyev:2022:QSV

[Afa22] G. A. Afanasyev. Queueing systems with vacations, interruptions, and delays in device operation. Theory of Probability and its Applications, 67(3):465–472, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasyev:2023:COQ

[Afa23] G. A. Afanasyev. Cost optimization of queueing systems with vacations. Theory of Probability and its Applications, 68(2):308–315, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ararat:2020:LCD

[AH20]  Ç. Ararat and A. H. Hamel. Lower

cone distribution functions and set-valued quantiles form Galois connections. Theory of Probability and its Applications, 65 (2):179–190, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Agram:2022:OSR

[AHØP22] N. Agram, S. Haadem, B. Øksendal, and F. Proske. Optimal stopping, randomized stopping, and singular control with general information flow. Theory of Probability and its Applications, 66(4):601–612, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T990642.

Aurzada:2022:APE

[AK22] F. Aurzada and M. Kilian.

Asymptotics of the persistence exponent of integrated fractional Brownian motion and fractionally integrated Brownian motion. Theory of Probability and its Applications, 67(1):77–88, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T990769.

Alexeev:2022:SRV

[Ale22] I. A. Alexeev. Stable random variables with complex stability index, I. Theory of Probability and its Applications, 67(3):335–351, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alexeev:2023:SRV

[Ale23] I. A. Alexeev. Stable random variables with complex stability index, II. Theory of Probability and its Applications, 67(4):499–515, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alimov:2020:MVR

[Ali20] A. R. Alimov. In memory of V.

R. Fatalov. Theory of Probability and its Applications, 64(4):656, ???? 2020. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aparin:2022:STD

[APY22] A. A. Aparin, G. A. Popov, and E. B. Yarovaya. On the sojourn time distribution of a random walk at a multidimensional lattice point. Theory of Probability and its Applications, 66(4):522–536, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

                                        /epubs.siam.org/doi/10.1137/
                                        S0040585X97T990599.

Arkin:2020:OTS

[Ark20] V. I. Arkin. Optimality of threshold stopping times for diffusion processes. Theory of Probability and its Applications, 65(3): 341–358, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bakay:2021:LDT

[Bak21] G. A. Bakay. Large deviations for a terminating compound renewal process. Theory of Probability and its Applications, 66 (2):209–227, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Balashova:2022:CEM

[Bal22] D. M. Balashova. Clustering effect for multitype branching random walk. Theory of Probability and its Applications, 67 (3):352–362, ???? 2022. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Biret:2022:GCT

[BBC22] M. Biret, M. Broniatowski, and Z. Cao. A Gibbs conditional theorem under extreme deviation. Theory of Probability and its Applications, 67(3):389–414, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belfadli:2023:SGM

[BCE23] R. Belfadli, M. Chadad, and M. Erraoui. On the sum of Gaussian martingale and an independent fractional Brownian motion. Theory of Probability and its Applications, 68(2):316–323, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Buckdahn:2022:MDH

[BDS22] R. Buckdahn, P. Di Tella, and W. Schmidt Shiryaev. In memoriam: Doctor Hans-Jürgen Engelbert. Theory of Probability and its Applications, 67(1):161–162, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T990848.

Belopolskaya:2021:SNB

[Bel21] Ya. I. Belopol’skaya. Systems of nonlinear backward and forward Kolmogorov equations: Generalized solutions. Theory of Probability and its Applications, 66 (1):15–43, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bingham:2022:LWL

[Bin22] N. H. Bingham. The life, work, and legacy of P. L. Chebyshev. Theory of Probability and its Applications, 66(4):506–521, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T990587.

Borovkov:2022:CTI

[BLM22] A. A. Borovkov, A. V. Logachov, and A. A. Mogulskii. Chebyshev-type inequalities and large deviation principles. Theory of Probability and its Applications, 66(4):570–581, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T990629.

Bobkov:2020:LLT

[BM20] S. G. Bobkov and A. Marsiglietti. Local limit theorems for smoothed Bernoulli and other convolutions. Theory of Probability and its Applications, 65(1):62–81, ???? 2020. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2020:PDR

[BMP20] A. A. Borovkov, A. A. Mogulskii, and E. I. Prokopenko. Properties of the deviation rate function and the asymptotics for the Laplace transform of the distribution of a compound renewal process. Theory of Probability and its Applications, 64(4):499–512, ???? 2020.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Babilua:2022:NEM

[BN22] P. Babilua and E. A. Nadaraya.

On a new estimation method of the Bernoulli regression function. Theory of Probability and its Applications, 67(2):163–174, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T99085X.

Bogachev:2022:CHP

[Bog22] V. I. Bogachev. Chebyshev– Hermite polynomials and distributions of polynomials in Gaussian random variables. Theory of Probability and its Applications, 66(4):550–569, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T990617.

Boldin:2020:PCS

[Bol20] M. V. Boldin. On the Pearson’s chi-square test for normality of autoregression with outliers. Theory of Probability and its Applications, 65(1):102–110, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2021:EIP

[Bor21a] A. A. Borovkov. Extension of the invariance principle for compound renewal processes to the zones of moderately large and small deviations. Theory of Probability and its Applications, 65 (4):511–526, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2021:ELD

[Bor21b] A. A. Borovkov. On exact large deviation principles for compound renewal processes. Theory of Probability and its Applications, 66(2):170–183, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Breton:2020:IRF

[BP20] J.-Ch. Breton and N. Privault. Integrability and regularity of the flow of stochastic differential equations with jumps. Theory of Probability and its Applications, 65 (1):82–101, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bruss:2022:GWP

[Bru22] F. Thomas Bruss. Galton–Watson processes and their role as building blocks for branching processes. Theory of Probability and its Applications, 67(1):141–153, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

                                        /epubs.siam.org/doi/10.1137/
                                        S0040585X97T990812.

Balint:2020:LFM

[BS20] D. Á. Bálint and M. Schweizer. Large financial markets, discounting, and no asymptotic arbitrage. Theory of Probability and its Applications, 65(2):191–223, ???? 2020. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borisov:2023:CRA

[BS23] I. S. Borisov and E. I. Shefer. Convergence rates to the arcsine law. Theory of Probability and its Applications, 68 (2):175–197, ???? 2023. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Banerjee:2023:GRO

[BSZ23] C. Banerjee, L. A. Sakhanenko, and D. C. Zhu. Global rate optimality of integral curve estimators in high order tensor models. Theory of Probability and its Applications, 68(2):250–266, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Barrasso:2022:CDM

[BT22] A. Barrasso and N. Touzi. Controlled diffusion mean field games with common noise and McKean– Vlasov second order backward SDEs. Theory of Probability and its Applications, 66(4):613–639, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T990654.

Bobkov:2022:CEC

[BU22] S. G. Bobkov and V. V. Ulyanov.

The Chebyshev–Edgeworth correction in the Central Limit Theorem for integer-valued independent summands. Theory of Probability and its Applications, 66(4):537–549, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

                                        epubs.siam.org/doi/10.1137/
                                        S0040585X97T990605.

Bulinskaya:2020:FPF

[Bul20] E. Vl. Bulinskaya. Fluctuations of the propagation front of a catalytic branching walk. Theory of Probability and its Applications, 64(4):513–534, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chebotarev:2023:SVN

[CBTV23] V. I. Chebotarev, I. S. Borisov, V. A. Topchii, and V. A. Vatutin. Sergey Viktorovich Nagaev (on his 90th birthday). Theory of Probability and its Applications, 68 (2):305–307, ???? 2023. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Choulli:2020:SCU

[CD20] T. Choulli and J. Deng. Structure conditions under progressively added information. Theory of Probability and its Applications, 65(3):418–453, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:2022:AFC

[CG22] G. P. Chistyakov and F. Götze. Approximation of free convolutions by free infinitely divisible laws. Theory of Probability and its Applications, 66(4):640–667, ???? 2022. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T990666.

Cuchiero:2020:FTA

[CKT20] C. Cuchiero, I. Klein, and J. Teichmann. A fundamental theorem of asset pricing for continuous time large financial markets in a two filtration setting. Theory of Probability and its Applications, 65(3):388–404, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cvitanic:2020:ICS

[CPRS20] J. Cvitanić, D. Prelec, S. Radas, and H. Sikić. Incentive-compatible surveys via posterior probabilities. Theory of Probability and its Applications, 65(2):292–321, ???? 2020. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chau:2020:BIC

[CR20] H. N. Chau and M. Rásonyi. Behavioral investors in conic market models. Theory of Probability and its Applications, 65 (2):330–337, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Choulli:2022:LOP

[CY22] T. Choulli and S. Yansori. Logoptimal portfolio without NFLVR: Existence, complete characterization, and duality. Theory of Probability and its Applications, 67(2):229–245, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

                                        epubs.siam.org/doi/10.1137/
                                        S0040585X97T990897.

Darkhovsky:2020:CDC

[Dar20] B. S. Darkhovsky. On the complexity and dimension of continuous finite-dimensional maps. Theory of Probability and its Applications, 65(3):375–387, ???? 2020. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dzindzalieta:2020:HSI

[DG20] D. Dzindzalieta and F. Götze. Half-spaces with influential variable. Theory of Probability and its Applications, 65(1):114–120, ???? 2020. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Davydov:2021:RAI

[DN21] Y. A. Davydov and S. M. Novikov. Remarks on asymptotic independence. Theory of Probability and its Applications, 66(1):44–57, ???? 2021. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dorpinghaus:2023:OIU

[DNRJ23] M. Dörpinghaus, I. Neri, É. Roldán, and F. Jülicher. Optimal information usage in binary sequential hypothesis testing. Theory of Probability and its Applications, 68(1):77–87, ???? 2023. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

deChaumaray:2020:SLD

[dRdC20] M. du Roy de Chaumaray. Sharp large deviations for the drift parameter of the explosive Cox– Ingersoll–Ross process. Theory of Probability and its Applications, 65(3):454–469, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Demidovich:2022:TLV

[DS22] Yu. A. Demidovich and D. A.

Shabanov. On two limit values of the chromatic number of a random hypergraph. Theory of Probability and its Applications, 67(2):175–193, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T990861.

EBTPA:2020:MDM

[Edi20a] Editorial Board of TPA. In memory of D. M. Chibisov (12.25.1936– 11.20.2019). Theory of Probability and its Applications, 65 (1):173–174, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

EBTPA:2020:MVM

[Edi20b] Editorial Board of TPA. In memory of V. M. Zolotarev (02.27.1931–11.07.2019). Theory of Probability and its Applications, 65(1):175–178, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

EBTPA:2020:CSK

[Edi20c] Editorial Board of TPA. On the centenary of Sagdy Khasanovich Sirazhdinov. Theory of Probability and its Applications, 65 (2):338–339, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

EBTPA:2021:BN

[Edi21a] Editorial Board of TPA. On the 75th birthday of A. A. Novikov. Theory of Probability and its Applications, 65(4):639, ???? 2021. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

EBTPA:2021:BB

[Edi21b] Editorial Board of TPA. On the 90th birthday of A. A. Borovkov. Theory of Probability and its Applications, 66(2):169, ???? 2021. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

EBTPA:2023:MGA

[Edi23] Editorial Board of TPA. In memory of L. G. Afanasyeva (14.09.1937–26.08.2022). Theory of Probability and its Applications, 67(4):668–670, ???? 2023. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

EsquIvel:2021:SMC

[EMP21] M. L. EsquÍvel, P. P. Mota, and J. P. Pina. On a stochastic model for a cooperative banking scheme for microcredit. Theory of Probability and its Applications, 66 (2):326–335, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Edelsbrunner:2020:WPD

[EN20] H. Edelsbrunner and A. Nikitenko.

Weighted Poisson–Delaunay mosaics. Theory of Probability and its Applications, 64(4):595–614, ???? 2020. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ferreira:2020:TDU

[FF20] M. Ferreira and H. Ferreira. Tail dependence under sample failures. Theory of Probability and its Applications, 64(4):636–645, ???? 2020. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feinberg:2020:FLW

[FKL20a] E. A. Feinberg, P. O. Kasyanov, and Y. Liang. Fatou’s lemma for weakly converging measures under the uniform integrability condition. Theory of Probability and its Applications, 64(4):615–630, ???? 2020. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feinberg:2020:FLC

[FKL20b] E. A. Feinberg, P. O. Kasyanov, and Y. Liang. Fatou’s lemma in its classical form and Lebesgue’s convergence theorems for varying measures with applications to Markov decision processes. Theory of Probability and its Applications, 65(2):270–291, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Formanov:2022:APS

[For22] Sh. K. Formanov. Another proof of a Sakhanenko theorem. Theory of Probability and its Applications, 67(3):473–477, ???? 2022. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feinberg:2022:KEJ

[FS22] E. A. Feinberg and A. N. Shiryaev. Kolmogorov’s equations for jump Markov processes and their applications to control problems. Theory of Probability and its Applications, 66(4):582–600, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T990630.

Giuliano:2020:SGC

[GCV20] R. Giuliano, M. Ordóñez Cabrera, and A. Volodin. On the sub-Gaussianity of the r-correlograms. Theory of Probability and its Applications, 65(3):470–481, ???? 2020. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Griffiths:2021:MDM

[GH21] R. Griffiths and K. Hamza. Matching the distributions of the marginals and the sums for the Meixner class. Theory of Probability and its Applications, 66 (3):430–444, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Giraudo:2021:EIS

[Gir21] D. Giraudo. An exponential inequality for U-statistics of I.I.D. data. Theory of Probability and its Applications, 66(3):408–429, ???? 2021. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Goto:2022:LRP

[GKKT22] Y. Goto, T. Kaneko, S. Kojima, and M. Taniguchi. Likelihood ratio processes under nonstandard settings. Theory of Probability and its Applications, 67(2):246–260, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T990903.

Gliklikh:2021:SSD

[Gli21] Yu. E. Gliklikh. On solvability of stochastic differential equations with osmotic velocities. Theory of Probability and its Applications, 65(4):640–647, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gotze:2020:MIL

[GNT20] F. Götze, A. A. Naumov, and A. N. Tikhomirov. Moment inequalities for linear and nonlinear statistics. Theory of Probability and its Applications, 65(1):1–16, ???? 2020. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grandits:2020:RPT

[Gra20] P. Grandits. A ruin problem for a two-dimensional Brownian motion with controllable drift in the positive quadrant. Theory of Probability and its Applications, 64 (4):646–655, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gushchin:2021:JLM

[Gus21] A. A. Gushchin. The joint law of a max-continuous local submartingale and its maximum. Theory of Probability and its Applications, 65(4):545–557, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gotze:2022:AAD

[GZ22] F. Götze and A. Yu. Zaitsev. On alternative approximating distributions in the multivariate version of Kolmogorov’s Second Uniform Limit Theorem. Theory of Probability and its Applications, 67(1):1–16, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T99071X.

Harlamov:2022:DDF

[Har22] B. P. Harlamov. Distribution density of the first exit point of a two-dimensional diffusion process from a circle neighborhood of its initial point: The inhomogeneous case. Theory of Probability and its Applications, 67(2):194–207, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T990873.

He:2020:LTR

[HB20] P. He and K. A. Borovkov. Limit theorems for record indicators in threshold F^{α}-schemes. Theory of Probability and its Applications, 65(3):405–417, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Iasnogorodski:2023:MVM

[ILM^{+}23] R. M. Iasnogorodski, A. A.

Lykov, M. V. Melikian, S. A. Pirogov, T. S. Turova-Shmeling, and A. A. Zamyatin. In memory of V. A. Malyshev (04.13.1938– 09.30.2022). Theory of Probability and its Applications, 68 (1):170–173, ???? 2023. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:2021:MLR

[IPF21] I. A. Ibragimov, E. L. Presman, and Sh. K. Formanov. On modifications of the Lindeberg and Rotar’ conditions in the Central Limit Theorem. Theory of Probability and its Applications, 65 (4):648–651, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:2021:AWP

[ISF21] I. A. Ibragimov, N. V. Smorodina, and M. M. Faddeev. Approximation of a Wiener process local time by functionals of random walks. Theory of Probability and its Applications, 66(1):58–74, ???? 2021. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:2022:RLP

[ISF22] I. A. Ibragimov, N. V. Smorodina, and M. M. Faddeev. Reflecting Lévy processes and associated families of linear operators. II. Theory of Probability and its Applications, 67(1):17–27, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

                                        /epubs.siam.org/doi/10.1137/
                                        S0040585X97T990721.

Ivanov:2023:BEM

[Iva23] D. V. Ivanov. On the bounds for the expected maxima of random samples with known expected maxima of two samples of smaller size. Theory of Probability and its Applications, 68(1):2–15, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Janisch:2021:KSL

[Jan21] M. Janisch. Kolmogorov’s Strong Law of Large Numbers holds for pairwise uncorrelated random variables. Theory of Probability and its Applications, 66(2): 263–275, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jurek:2022:BDD

[Jur22] Z. J. Jurek. Background driving distribution functions and series representations for log-gamma self-decomposable random variables. Theory of Probability and its Applications, 67(1):105–117, ???? 2022. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T990782.

Kabanov:2021:SLC

[Kab21] Yu. M. Kabanov. On sets of laws of continuous martingales. Theory of Probability and its Applications, 65(4):652–655, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kaushik:2022:NSB

[Kau22] P. Kaushik. A new solution of Bertrand’s Paradox. Theory of Probability and its Applications, 67(1):158–160, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T990836.

Konakov:2021:CCC

[KF21] V. D. Konakov and A. R. Falaleev.

Convergence of certain classes of random flights in the Kantorovich metric. Theory of Probability and its Applications, 65(4):656–664, ???? 2021. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kharlamov:2022:ASS

[Kha22] V. V. Kharlamov. On asymptotic strategies in the stochastic Colonel Blotto game. Theory of Probability and its Applications, 67(2):318–326, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T990952.

Khatuntseva:2023:BI

[Kha23] M. V. Khatuntseva. On the 90th birthday of I. A. Ibragimov. Theory of Probability and its Applications, 68(1):1, ???? 2023. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khvorostyanskaya:2023:NTG

[Khv23] E. V. Khvorostyanskaya. On the number of trees of a given size in a Galton–Watson forest in the critical case. Theory of Probability and its Applications, 68(1):62–76, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klump:2023:UIF

[KK23] A. Klump and M. Kolb. Uniqueness of the inverse first-passage time problem and the shape of the Shiryaev boundary. Theory of Probability and its Applications, 67(4):570–592, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T991155.

Karagodin:2021:DLE

[KL21] N. A. Karagodin and M. A. Lifshits. On the distribution of the last exit time over a slowly growing linear boundary for a Gaussian process. Theory of Probability and its Applications, 66 (3):337–347, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korbel:2022:PFS

[KM22] S. Korbel and P. Mörters. A path formula for the sock sorting problem. Theory of Probability and its Applications, 66(4):708–712, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T99068X.

Kordzakhia:2022:MIO

[KN22] N. E. Kordzakhia and A. A.

Novikov. On maximal inequalities for Ornstein–Uhlenbeck processes with jumps. Theory of Probability and its Applications, 66(4):713–728, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T990691.

Kovacevic:2021:MES

[Kov21] M. Kovacević. On the maximum entropy of a sum of independent discrete random variables. Theory of Probability and its Applications, 66(3):482–487, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kogut:2021:TDD

[KR21] N. S. Kogut and I. V. Rodionov.

On tests for distinguishing distribution tails. Theory of Probability and its Applications, 66 (3):348–363, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kremena:2020:SHE

[Kre20] E. V. Kremena. On the shape of high excursions of Gaussian stationary processes. Theory of Probability and its Applications, 65 (1):111–113, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Karatzas:2022:TAG

[KST22] I. Karatzas, W. Schachermayer, and B. Tschiderer. A trajectorial approach to the gradient flow properties of Langevin– Smoluchowski diffusions. Theory of Probability and its Applications, 66(4):668–707, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T990678.

Klimov:2023:SWO

[KT23] K. Yu. Klimov and M. N. Tsareva. On student workshops organized by the Vega Institute. Theory of Probability and its Applications, 68(2):337–338, ???? 2023. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lang:2022:CMC

[LCY^{+}22] J. Lang, L. Cheng, Z. Yu, Y. Wu, and X. Wang. Complete fmoment convergence for randomly weighted sums of extended negatively dependent random variables and its statistical application. Theory of Probability and its Applications, 67(2):261–281, ???? 2022. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T990915.

Lebedev:2022:RIM

[Leb22] A. V. Lebedev. Records and increases of multivariate extremes of random particle scores in supercritical branching processes with max-linear heredity. Theory of Probability and its Applications, 67(2):310–317, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T990940.

Lee:2021:FPT

[Lee21] J. M. Lee. The first passage time density of Brownian motion and the heat equation with Dirichlet boundary condition in time dependent domains. Theory of Probability and its Applications, 66 (1):142–159, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lenena:2023:AEV

[Len23] E. O. Lenena. Accuracy of estimation of the vector of queue lengths for open Jackson networks. Theory of Probability and its Applications, 67(4):633–639, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T991180.

Lopez-Garcia:2020:IMP

[LG20] M. López-García. Indetermined moment problems related to qfunctional equations. Theory of Probability and its Applications, 65(3):482–496, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lv:2023:QSD

[LH23] Y. Lv and W. Hong. Quenched small deviation for the trajectory of a random walk with random environment in time. Theory of Probability and its Applications, 68(2):267–284, ???? 2023. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Li:2021:EEE

[Li21] Z. Li. Ergodicities and exponential ergodicities of Dawson–Watanabe type processes. Theory of Probability and its Applications, 66 (2):276–298, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Linke:2023:TIN

[Lin23] Yu. Yu. Linke. Towards insensitivity of Nadaraya–Watson estimators to design correlation. Theory of Probability and its Applications, 68(2):198–210, ???? 2023. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lifshits:2023:PPZ

[LL23] M. A. Lifshits and I. M. Lialinov. Probabilistic properties of Zipf sets and their maximal intersections. Theory of Probability and its Applications, 68(1):16–30, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lu:2023:SEF

[LMW23] T. Lu, C. Ma, and F. Wang.

Series expansions of fractional Brownian motions and strong local nondeterminism of bifractional Brownian motions on balls and spheres. Theory of Probability and its Applications, 68 (1):88–110, ???? 2023. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lifshits:2023:MSR

[LT23] M. A. Lifshits and A. A. Tadevosian. On the maximum of a special random assignment process. Theory of Probability and its Applications, 67(4):640–644, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T991192.

Lu:2021:CMC

[LWW21] C. Lu, X. J. Wang, and Y. Wu.

Complete moment convergence for the dependent linear processes with application to the state observers of linear-time-invariant systems. Theory of Probability and its Applications, 65(4):570–587, ???? 2021. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lu:2022:NBE

[LYJ^{+}22] C. Lu, W. Yu, R. L. Ji, H. L.

Zhou, and X. J. Wang. A note on the Berry–Esseen bounds for ρmixing random variables and their applications. Theory of Probability and its Applications, 67 (3):415–433, ???? 2022. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mousavinasr:2022:MDC

[MGD22] S. Mousavinasr, C. R. Gonçalves, and C. C. Y. Dorea. Mallows distance convergence for extremes: Regeneration approach. Theory of Probability and its Applications, 67(3):478–484, ???? 2022. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mosyagin:2021:ADT

[Mos21] V. E. Mosyagin. Asymptotics for the distribution of the time of attaining the maximum for a trajectory of a Poisson process with linear drift and intensity switch. Theory of Probability and its Applications, 66(1):75–88, ???? 2021. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Murzintseva:2023:HPA

[MPP23] A. A. Murzintseva, S. M. Pergamenchtchikov, and E. A. Pchelintsev. Hedging problem for Asian call options with transaction costs. Theory of Probability and its Applications, 68 (2):211–230, ???? 2023. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Miao:2022:EGA

[MQY22] Y. Miao, X. M. Qu, and G. Y.

Yang. Extrema of the generalized allocation scheme based on an m-dependent sequence. Theory of Probability and its Applications, 67(2):282–293, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T990927.

Melkumova:2021:CQR

[MS21] L. E. Melkumova and S. Ya.

Shatskikh. Conditional quantile reproducibility of multivariate distributions and simplified pair copula construction. Theory of Probability and its Applications, 66 (1):160–168, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maller:2023:GPD

[MS23] R. Maller and S. Shemehsavar.

Generalized Poisson–Dirichlet distributions based on the Dickman subordinator. Theory of Probability and its Applications, 67(4):593–612, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T991167.

Nagaev:2021:AMP

[Nag21a] S. V. Nagaev. An alternative method of the proof of the ergodic theorem for general Markov chains. Theory of Probability and its Applications, 66(3):364–375, ???? 2021. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:2021:ESS

[Nag21b] S. V. Nagaev. An estimate for the sum of the Spitzer series and its generalization. Theory of Probability and its Applications, 66 (1):89–104, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nguyen:2020:AHC

[NP20] T. Nguyen and S. Pergamenschchikov. Approximate hedging with constant proportional transaction costs in financial markets with jumps. Theory of Probability and its Applications, 65 (2):224–248, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nguyen:2021:MTH

[NP21] T. Nguyen and H. Pham. A maximal theorem of Hardy–Littlewood type for pairwise I.I.D. random variables and the Law of Large Numbers. Theory of Probability and its Applications, 66(3): 445–454, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

OrdonezCabrera:2022:NVU

[ORÜV22] M. Ordóñez Cabrera, A. Rosalsky, M. Ünver, and A. Volodin. A new version of uniform integrability via power series summability methods. Theory of Probability and its Applications, 67(1):89–104, ???? 2022. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T990770.

ORourke:2023:PLE

[OW23] S. O’Rourke and N. Williams.

Partial linear eigenvalue statistics for non-Hermitian random matrices. Theory of Probability and its Applications, 67(4):613–632, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T991179.

Palamarchuk:2020:UFI

[Pal20] E. S. Palamarchuk. On upper functions for integral quadratic functionals based on time-varying Ornstein–Uhlenbeck process. Theory of Probability and its Applications, 65(1):17–31, ???? 2020. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Palamarchuk:2022:OSL

[Pal22] E. S. Palamarchuk. On optimal stochastic linear quadratic control with inversely proportional time-weighting in the cost. Theory of Probability and its Applications, 67(1):28–43, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T990733.

Palamarchuk:2023:OLR

[Pal23a] E. S. Palamarchuk. On optimal linear regulator with polynomial process of external excitations. Theory of Probability and its Applications, 67(4):535–547, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

                                        /epubs.siam.org/doi/10.1137/
                                        S0040585X97T991131.

Palamarchuk:2023:OLQ

[Pal23b] E. S. Palamarchuk. Optimal linear-quadratic regulator for a stochastic system under mutually inverse time preferences in the cost. Theory of Probability and its Applications, 68(1):31–45, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pavlov:2021:LDN

[PC21] Yu. L. Pavlov and I. A. Cheplyukova. Limit distributions of the number of vertices of a given degree in a configuration graph with bounded number of edges. Theory of Probability and its Applications, 66(3):376–390, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pinelis:2022:ELU

[Pin22] I. Pinelis. Exact lower and upper bounds for Gaussian measures. Theory of Probability and its Applications, 67(3):485–493, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pinelis:2023:ARE

[Pin23] I. Pinelis. Asymptotic relative efficiency of the Kendall and Spearman correlation statistics. Theory of Probability and its Applications, 68(1):111–122, ???? 2023. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Piunovskiy:2023:TFM

[Piu23] A. B. Piunovskiy. Turnpikes in finite Markov decision processes and random walk. Theory of Probability and its Applications, 68 (1):123–149, ???? 2023. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Platonova:2022:AFK

[Pla22]  M. V. Platonova.  An ana-

logue of the Feynman–Kac formula for a high-order operator. Theory of Probability and its Applications, 67(1):62–76, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T990757.

Platonova:2023:MAP

[PR23] M. V. Platonova and K. S.

Ryadovkin. Moment asymptotics of population size of particles at vertices for a supercritical branching random walk on a periodic graph. Theory of Probability and its Applications, 68(2):231–249, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pawlas:2021:RMN

[PS21a] P. Pawlas and D. Szynal. On relations for moments of the normal k-th record values and a characterization. Theory of Probability and its Applications, 66 (3):488–494, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:2021:CST

[PS21b] A. V. Prokhorov and A. V. Savitskii. Characterizing statistics and their applications. Theory of Probability and its Applications, 66(3):391–407, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Piterbarg:2022:AMA

[PS22] V. I. Piterbarg and Yu. A. Scherbakova. On accompanying measures and asymptotic expansions in the B. V. Gnedenko Limit Theorem. Theory of Probability and its Applications, 67(1):44–61, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

                                        /epubs.siam.org/doi/10.1137/
                                        S0040585X97T990745.

Platonova:2021:PAS

[PT21] M. V. Platonova and S. V. Tsykin. Probabilistic approximation of the solution of the Cauchy problem for the higher-order Schrödinger equation. Theory of Probability and its Applications, 65(4): 558–569, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Perfilev:2022:LTA

[PW22] E. Perfilev and V. Wachtel. Local tail asymptotics for the joint distribution of the length and of the maximum of a random walk excursion. Theory of Probability and its Applications, 67(2):294–309, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T990939.

Peng:2022:DUF

[PY22] S. Peng and S. Yang. Distributional uncertainty of the financial time series measured by G-expectation. Theory of Probability and its Applications, 66(4):729–741, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T990708.

Rokhlin:2021:RAC

[Rok21] D. B. Rokhlin. Resource allocation in communication networks with large number of users: The dual stochastic gradient method. Theory of Probability and its Applications, 66 (1):105–120, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Roos:2022:ACC

[Roo22] B. Roos. On the accuracy in a combinatorial Central Limit Theorem: The characteristic function method. Theory of Probability and its Applications, 67(1):118–139, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T990794.

Rozovsky:2020:ILC

[Roz20] L. V. Rozovsky. On integro-local CLT for sums of independent random vectors. Theory of Probability and its Applications, 64 (4):564–578, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovsky:2022:LDS

[Roz22] L. V. Rozovsky. Large deviations of a sum of independent random variables with distributions with rapidly decreasing tails. Theory of Probability and its Applications, 67(3):363–374, ???? 2022. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovsky:2023:CRP

[Roz23] L. V. Rozovsky. On the convergence rate in precise asymptotics. Theory of Probability and its Applications, 68(1):46–61, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rubinshtein:2021:PBM

[RS21] A. I. Rubinshtein and V. B. Sherstyukov. On a probabilistic Bernstein model. Theory of Probability and its Applications, 66 (2):318–325, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Savitskii:2020:LTF

[Sav20] A. V. Savitskii. Limit theorems for functions of a fractional Brownian motion. Theory of Probability and its Applications, 65(1):32–48, ???? 2020. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Savelov:2021:TSC

[Sav21] M. P. Savelov. Two-stage chisquare test and two-dimensional distributions of a Bessel process. Theory of Probability and its Applications, 65(4):665–672, ???? 2021. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Son:2022:GML

[SDDT22] T. C. Son, L. V. Dung, D. T.

Dat, and T. T. Trang. Generalized Marcinkiewicz laws for weighted dependent random vectors in Hilbert spaces. Theory of Probability and its Applications, 67(3):434–451, ???? 2022. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2021:MVV

[Shi21] A. N. Shiryaev. In memory of V. V. Senatov (03.10.1951– 06.22.2021). Theory of Probability and its Applications, 66 (3):495–496, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2022:BT

[Shi22a] A. N. Shiryaev. Birthday tributes. Theory of Probability and its Applications, 67(1):140, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

/epubs.siam.org/doi/10.1137/
S0040585X97T990800.

Shiryaev:2022:BBP

[Shi22b] A. N. Shiryaev. On the bicentenary of the birth of P. L. Chebyshev, a great Russian mathematician. Theory of Probability and its Applications, 66(4):497–505, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https:/

                                        /epubs.siam.org/doi/10.1137/
                                        S0040585X97T990575.

Sobolev:2022:SMA

[SK22] V. N. Sobolev and A. E. Kondratenko. On Senatov moments in asymptotic expansions in the Central Limit Theorem. Theory of Probability and its Applications, 67(1):154–157, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

                                        epubs.siam.org/doi/10.1137/
                                        S0040585X97T990824.

Skorski:2023:SGC

[Sko23] M. Skorski. On sub-Gaussian concentration of missing mass. Theory of Probability and its Applications, 68(2):324–329, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Slepov:2021:CRR

[Sle21] N. A. Slepov. Convergence rate of random geometric sum distributions to the Laplace law. Theory of Probability and its Applications, 66(1):121–141, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Stoyanov:2020:NCC

[SLK20] J. M. Stoyanov, G. D. Lin, and P. Kopanov. New checkable conditions for moment determinacy of probability distributions. Theory of Probability and its Applications, 65(3):497–509, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Smirnov:2023:GDA

[Smi23] S. N. Smirnov. A guaranteed deterministic approach to superhedging: The relationship between the deterministic and probabilistic problem statements without trading constraints. Theory of Probability and its Applications, 67(4):548–569, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T991143.

Shiryaev:2020:ATG

[SP20] A. N. Shiryaev and I. V. Pavlov.

Abstracts of talks given at the 4th International Conference on Stochastic Methods. Theory of Probability and its Applications, 65(1):121–172, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2023:ATG

[SP23a] A. N. Shiryaev and I. V. Pavlov.

Abstracts of talks given at the 7th International Conference on Stochastic Methods, I. Theory of Probability and its Applications, 67(4):652–667, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

                                        epubs.siam.org/doi/10.1137/
                                        S0040585X97T991210.

Shiryaev:2023:ATGb

[SP23b] A. N. Shiryaev and I. V. Pavlov. Abstracts of talks given at the 7th International Conference on Stochastic Methods, II. Theory of Probability and its Applications, 68(1):150–169, ???? 2023. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Salimov:2020:AMS

[SS20] R. F. Salimov and S. V. Simushkin. Asymptotics of the minimum sufficient number of observations for d-guaranteed discrimination of two-sided hypotheses. Theory of Probability and its Applications, 65(1):49–61, ???? 2020. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Spielmann:2020:RPI

[SV20] J. Spielmann and L. Vostrikova. On the ruin problem with investment when the risky asset is a semimartingale. Theory of Probability and its Applications, 65 (2):249–269, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Soshnikov:2021:GAD

[SX21] A. Soshnikov and Y. Xu. Gaussian approximation of the distribution of strongly repelling particles on the unit circle. Theory of Probability and its Applications, 65(4):588–615, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sun:2021:SAP

[SZ21] S. Sun and W. Zhu. Some asymptotic properties between smooth empirical and quantile processes for dependent random variables. Theory of Probability and its Applications, 66(3): 455–468, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Timashev:2021:LLT

[Tim21] A. N. Timashev. Local limit theorems for compound discrete distributions. Theory of Probability and its Applications, 66 (3):474–481, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Toofanpour:2022:NLL

[TJIN22] J. Toofanpour, M. Javanian, and R. Imany-Nabiyyi. Normal limit law for protected node profile of random recursive trees. Theory of Probability and its Applications, 67(3):452–464, ???? 2022. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tu:2021:AOF

[Tu21] E. Tu. Association and other forms of positive dependence for Feller evolution systems. Theory of Probability and its Applications, 66(2):299–317, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:2020:IES

[VD20] V. A. Vatutin and E. E. Dyakonova. The initial evolution stage of a weakly subcritical branching process in a random environment. Theory of Probability and its Applications, 64 (4):535–552, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:2021:SBP

[VD21] V. A. Vatutin and E. E. Dyakonova. Subcritical branching processes in random environment with immigration: Survival of a single family. Theory of Probability and its Applications, 65 (4):527–544, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:2023:APS

[VD23] V. A. Vatutin and E. E. D’yakonova. Atypical population size in a two-type decomposable branching process. Theory of Probability and its Applications, 67(4):516–534, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T99112X.

Veretennikov:2021:PUS

[Ver21] A. Yu. Veretennikov. On pathwise uniqueness of solutions for multidimensional McKean–Vlasov equation. Theory of Probability and its Applications, 66(3):

469–473, ????  2021.  CO-

DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Xi:2023:CCI

[XLCW23] M. M. Xi, X. Q. Li, L. Chen, and X. J. Wang. Complete and complete integral convergence for arrays of rowwise extended negatively dependent random variables under sublinear expectations. Theory of Probability and its Applications, 68(2):285–304, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yarovaya:2022:LE

[Yar22a] E. B. Yarovaya. Letter to the Editors. Theory of Probability and its Applications, 67(3):498, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yarovaya:2022:NSL

[Yar22b] E. B. Yarovaya. News of scientific life — information on the General Seminar of the Department of Probability, Faculty of Mathematics and Mechanics, Lomonosov Moscow State University, Moscow, Russia, 2021. Theory of Probability and its Applications, 67(2):327–334, ???? 2022. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL https://

epubs.siam.org/doi/10.1137/
S0040585X97T990964.

Yarovaya:2022:NSLb

[Yar22c] E. B. Yarovaya. News of scientific life — information on the General Seminar of the Department of Probability, Faculty of Mechanics and Mathematics, Lomonosov Moscow State University, Moscow, Russia, Spring Term 2022. Theory of Probability and its Applications, 67 (3):494–497, ???? 2022. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yarovaya:2023:IGS

[Yar23] E. B. Yarovaya. Information on the General Seminar of the Department of Probability, Faculty of Mechanics and Mathematics, Lomonosov Moscow State University, Moscow, Russia, Fall Term 2022. Theory of Probability and its Applications, 68(2):330–333, ???? 2023. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yaskov:2022:LSD

[Yas22] P. A. Yaskov. Limiting spectral distribution for large sample covariance matrices with graphdependent elements. Theory of Probability and its Applications, 67(3):375–388, ???? 2022. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yarovaya:2020:CPD

[YSK20] E. B. Yarovaya, J. M. Stoyanov, and K. K. Kostyashin. On conditions for a probability distribution to be uniquely determined by its moments. Theory of Probability and its Applications, 64 (4):579–594, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zadorozhniy:2021:ESL

[Zad21] V. G. Zadorozhniy. The expectation of a solution of a linear system of differential equations with random coefficients. Theory of Probability and its Applications, 66(2):228–244, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhitlukhin:2020:SPS

[Zhi20]  M. V. Zhitlukhin.  Support-

ing prices in a stochastic von Neumann–Gale model of a financial market. Theory of Probability and its Applications, 64 (4):553–563, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhitlukhin:2023:SCB

[Zhi23] M. V. Zhitlukhin. Seminars, conferences, books. Theory of Probability and its Applications, 68 (2):334–336, ???? 2023. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhao:2020:CGT

[ZL20] J. Zhao and E. Lépinette. A complement to the Grigoriev theorem for the Kabanov model. Theory of Probability and its Applications, 65(2):322–329, ???? 2020. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zorich:2020:RDV

[Zor20] V. A. Zorich. A remark on the distribution of values for functions of a large number of variables. Theory of Probability and its Applications, 64(4):631–635, ???? 2020. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhou:2021:PDR

[ZSG21] Q. Zhou, A. I. Sakhanenko, and J. Guo. Prokhorov distance with rates of convergence under sublinear expectations. Theory of Probability and its Applications, 65(4):616–638, ???? 2021. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).